1. Below is a function f for which f and t” have already been computed for you. f(x) = 24 – 43% + 162 (t) = 4(x + 1)(x - 2

Please answer clearly and step by step, thank you!!!!
Answer #1

@ 7 foxis 24-4n3 +164, flx) = 4 (x+4) (x-ze; f(X) = 12x12-2) for Incoreasing f(x180 > 41X+0(x-22 yo (X+1) 70 (x+ (-1,00)]Tfixa 4x + 25 -3 XG [11253 Flixla4-25 o to find critical points X2 ux-1=0 *= 25 - 52 Fon= 4+28-3=26 425)= 100 +-3 = 98 H5727=